# Optimal Computation in Anonymous Dynamic Networks

### v7

@@ 5.3 Subroutine ApproxCount
We will now define the subroutine $\texttt{ApproxCount}(\mathcal{V},s,x,\ell)$ introduced in Section 5.2 and invoked in Listing 2. We will also give a proof to the technical Lemma 5.1 (restated as Lemma 5.9).
Overview. In ApproxCount, we are given a view $\mathcal{V}$ with a strand of leader nodes hanging from the first leader node $\vartheta$ in level $L_{s}$, where the anonymity $a(\vartheta)$ is an unknown number not greater than $\ell$. The algorithm begins by assuming that $a(\vartheta)$ is the given parameter $x$, and then it makes deductions on the anonymities of other nodes until it is able to make an estimate $n^{\prime}>0$ on the total number of processes, or report failure in the form of an error code $n^{\prime}\in\{-1,-2,-3\}$. In particular, since the algorithm requires the existence of a long-enough strand hanging from $\vartheta$, it reports failure if some descendants of $\vartheta$ (in the relevant levels of $\mathcal{V}$) have more than one child.
An important difficulty that is unique to the multi-leader case is that, even if $\mathcal{V}$ contains a long-enough strand of leader nodes, some nodes in the strand may still be branching in the history tree, and all branches except one may be missing from $\mathcal{V}$.
@@
Discrepancy $\delta$. Suppose that ApproxCount is invoked with arguments $\mathcal{V}$, $s$, $x$, $\ell$, where $1\leq x\leq\ell$, and let $\vartheta$ be the first leader node in level $L_{s}$ of $\mathcal{V}$ (if $\vartheta$ does not exist, the procedure immediately returns the error code $n^{\prime}=-1$). We define the discrepancy $\delta$ as the ratio $x/a(\vartheta)$. Clearly, $1/\ell\leq\delta\leq\ell$. Note that, since $a(\vartheta)$ is not a-priori known by the process executing ApproxCount, then neither is $\delta$.
@@
Let $u,v^{\prime}\in L_{t}$, and let $P_{1}$ and $P_{2}$ be the sets of processes represented by $u$ and $v^{\prime}$, respectively. By counting the links between $P_{1}$ and $P_{2}$ in $G_{t+1}$ in two ways, we have
@@
Our proof is by well-founded induction on $w(v)$. Assume for a contradiction that no descendants of $v$ are heavy. Let $v_{1}$, $v_{2}$, …, $v_{k}$ be the “immediate” descendants of $v$ that have been assigned guesses. That is, for all $1\leq i\leq k$, no internal nodes of the black path with endpoints $v$ and $v_{i}$ have been assigned guesses. Observe that $k\geq 1$ because, by assumption, $w(v)>a^{\prime}(v)=\delta a(v)\geq a(v)\geq 1$.
@@
for all $1\leq i\leq k$. Also,
@@
Cuts and isles. Fix a view $\mathcal{V}$ of a history tree $\mathcal{H}$. A set of nodes $C$ in $\mathcal{V}$ is said to be a cut for a node $v\notin C$ of $\mathcal{V}$ if two conditions hold: (i) for every leaf $v^{\prime}$ of $\mathcal{V}$ that lies in the subtree hanging from $v$, the black path from $v$ to $v^{\prime}$ contains a node of $C$, and (ii) no proper subset of $C$ satisfies condition (i). A cut for the root $r$ whose nodes are all counted is said to be a counting cut.
@@
If $s$ is an isle’s root and $F$ is its set of leaves, we have $a(s)\geq\sum_{v\in F}a(v)$, which is equivalent to $a^{\prime}(s)\geq\sum_{v\in F}a^{\prime}(v)$. Note that equality does not necessarily hold, because $s$ may have some descendants in the history tree $\mathcal{H}$ that do not appear in the view $\mathcal{V}$. If equality holds, then the isle is said to be complete. In this case, given the conditional anonymities of $s$ and of all nodes in $F$, we can easily compute the conditional anonymities of all the internal nodes by adding them up starting from the nodes in $F$ and working our way up to $s$.
Figure 7: The first levels of (a view of) a history tree, where the colored nodes are counted. The blue nodes form a counting cut, and the orange ones define a non-trivial isle with root $s$, where the nodes with a dot are internal.
@@
1# Input: a view $\mathcal{V}$ and three integers $s$, $x$, $\ell$
2# Output: a pair of integers $(n^{\prime},t)$
@@ 32 Else return $(n^{\prime},t)$
Algorithmic details of ApproxCount. The algorithm ApproxCount uses flags to mark nodes as “guessed”, “counted”, or “locked”; initially, no node is marked. Thanks to these flags, we can check if a node $u\in\mathcal{V}$ is a guesser: let $u_{1}$, $u_{2}$, …, $u_{k}$ be the children of $u$ that are also in $\mathcal{V}$ (recall that a view does not contain all nodes of a history tree); $u$ is a guesser if and only if it is marked as counted, all the $u_{i}$’s are marked as counted, and $a^{\prime}(u)=\sum_{i}a^{\prime}(u_{i})$ (which implies $a(u)=\sum_{i}a(u_{i})$, and thus no children of $u$ are missing from $\mathcal{V}$).
@@
The algorithm starts by assigning a conditional anonymity $a^{\prime}(\vartheta)=x$ to the first leader node $\vartheta\in L_{s}$. If no leader node exists in $L_{s}$, the error code $-1$ is immediately returned (Line 6). The algorithm also finds the longest strand $P_{\vartheta}$ hanging from $\vartheta$, assigns the same conditional anonymity $x$ to all of its nodes (including the unique child of the last node of $P_{\vartheta}$) and marks them as counted (Lines 7–11). Then, as long as there are nodes that can receive a guess and no counting cut has been found, the algorithm keeps assigning guesses to nodes. A node can receive a guess if it is guessable, not counted, and not locked (Line 12).
When a guess is made on a node $v$, the node itself and all of its siblings become locked (Line 15). This is to ensure that guessed nodes will always be well spread. Moreover, as a result of $v$ becoming guessed, some nodes in the path from $v$ to its ancestor in $L_{s}$ may become heavy; if this happens, let $v^{\prime}$ be the deepest heavy node (Line 18). If $g(v^{\prime})$ is not an integer, the algorithm returns the error code $-3$ (Line 19). (As we will prove later, this can only happen if $\delta\neq 1$ or some nodes in the strand $P_{\vartheta}$ have children that are not in the view $\mathcal{V}$.) Otherwise, if $g(v^{\prime})$ is an integer, the algorithm marks $v^{\prime}$ as counted and not guessed, in accordance with Lemma 5.5. Also, since $v^{\prime}$ is no longer guessed, its siblings become unlocked and are again eligible to receive guesses (Line 20). Furthermore, if the newly counted node $v^{\prime}$ is either the root or a leaf of a complete isle $I$, then the conditional anonymities of all the internal nodes of $I$ are determined, and such nodes are marked as counted (Lines 21–24).
In the end, the algorithm performs a “reality check” and possibly returns an estimate $n^{\prime}$ of $n$ as follows. If no counting cut was found, the algorithm returns the error code $-2$ (Line 25). Otherwise, a counting cut $C$ has been found. The algorithm computes $n^{\prime}$ (respectively, $\ell^{\prime}$) as the sum of the conditional anonymities of all nodes (respectively, all leader nodes) in $C$. If $\ell^{\prime}=\ell$, then the algorithm returns $n^{\prime}$ (Line 32). Otherwise, it returns the error code $-1$ if $\ell^{\prime}<\ell$ (Line 30) or the error code $-3$ if $\ell^{\prime}>\ell$ (Line 31). In all cases, the algorithm also returns the maximum depth $t$ of a guessed or counted node (excluding $\vartheta$ and its descendants), or $s$ if no such node exists.
@@
Recall that a node $v$ of the history tree $\mathcal{H}$ is said to be missing from level $L_{i}$ of the view $\mathcal{V}$ if $v$ is at the level of $\mathcal{H}$ corresponding to $L_{i}$ but does not appear in $\mathcal{V}$. Clearly, if a level of $\mathcal{V}$ has no missing nodes, all previous levels have no missing nodes, either.
@@
Observe that the subtree $\mathcal{V}^{\prime}$ of $\mathcal{V}$ formed by all levels up to $L_{t}$ satisfies all of the above conditions, as it contains $\vartheta\in L_{s}$, the root $r$, and has no counting cuts, because a counting cut for $\mathcal{V}^{\prime}$ would be a counting cut for $\mathcal{V}$, as well (recall that $\mathcal{V}$ has no counting cuts). Also, Lemma 5.6 (iii) ensures that $\mathcal{V}^{\prime}$ contains no non-trivial complete isles. However, since no nodes are missing from the levels of $\mathcal{V}^{\prime}$, all isles in $\mathcal{V}^{\prime}$ are complete, and thus must be trivial. We conclude that, if $\mathcal{V}^{\prime}$ has $f$ leaves not in the subtree hanging from $\vartheta$, it contains at most $f-1$ bad nodes. Since such leaves induce a partition of the at most $n-1$ processes not represented by $\vartheta$, we have $f\leq n-1$, implying that the number of bad nodes up to $L_{t}$ is at most $n-2$. ∎
@@
    If $n^{\prime}>0$ and $t^{\prime}\geq t+n^{\prime}$ and $a(\vartheta)=a(\vartheta^{\prime})$, then $n^{\prime}=n$.
@@
    If $t^{\prime}\geq s+(\ell+2)(n-1)$, then $s\leq t\leq s+(\ell+1)(n-1)$ and $n^{\prime}\neq-1$. Moreover, if $n^{\prime}=-2$, then $L_{t}$ contains a leader node with at least two children in $\mathcal{V}$.
@@
Let us prove statement (i). The assumption $x=a(\vartheta)$ implies $\delta=1$. Moreover, since $a(\vartheta)=a(\vartheta^{\prime})$, 1 is satisfied whenever a guess is made (this is a straightforward induction). Therefore, by Lemma 5.6 (iv), all nodes marked as counted up to $L_{t}$ indeed have the correct guesses. So, the conditional anonymity that is computed for any node is equal to its anonymity ($a^{\prime}(v)=\delta a(v)=a(v)$), and hence is an integer. This implies that ApproxCount cannot return the error code $-3$ at Line 19. Also, either $\ell^{\prime}=\ell$ if all leader processes have been counted, or $\ell^{\prime}<\ell$ if some leader nodes are missing from the view. Either way, ApproxCount cannot return the error code $-3$ at Line 31. We conclude that $n^{\prime}\neq-3$.
Let us prove statement (ii). Again, because $a(\vartheta)=a(\vartheta^{\prime})$, 1 is satisfied, and all nodes marked as counted have correct guesses. Also, $x\geq a(\vartheta)$ is equivalent to $\delta\geq 1$. By assumption, ApproxCount returns $(n^{\prime},t)$ with $n^{\prime}>0$ and $t^{\prime}\geq t+n^{\prime}$. Since $n^{\prime}>0$, a counting cut $C$ was found whose nodes are within levels up to $L_{t}$, and $n^{\prime}$ is the sum of the conditional anonymities of all nodes in $C$. Let $S_{C}$ be the set of processes represented by the nodes of $C$; note that $n^{\prime}\geq|S_{C}|$, because $\delta\geq 1$. We will prove that $S_{C}$ includes all processes in the system. Assume the contrary; Lemma 3.3 implies that, since $t^{\prime}\geq t+n^{\prime}\geq t+|S_{C}|$, there is a node $z\in L_{t}$ representing some process not in $S_{C}$. Thus, the black path from $z$ to the root $r$ does not contain any node of $C$, contradicting the fact that $C$ is a counting cut with no nodes after $L_{t}$. Therefore, $|S_{C}|=n$, i.e., the nodes in $C$ represent all processes in the system. Since ApproxCount returns $n^{\prime}>0$, the “reality check” $\ell^{\prime}=\ell$ succeeds (Lines 30–32). However, $\ell^{\prime}$ is the sum of the conditional anonymities of all leader nodes in $C$, and hence $\ell^{\prime}=\delta\ell$, implying that $\delta=1$. Thus, $n^{\prime}=\delta n=n$, as claimed.
@@
Assume that $p<s+(\ell+1)(n-1)$. This implies that no nodes are missing from level $L_{p+1}$, and therefore $\vartheta^{\prime\prime}$ must have at least two children in $L_{p+1}$. Since $t\leq p$, we have $t<s+(\ell+1)(n-1)$, as desired. Now assume that $n^{\prime}=-2$, which implies that the algorithm was unable to find a counting cut. We claim that in this case $t=p$. So, assume for a contradiction that $t\leq p-1$. It follows that $\vartheta^{\prime}\in L_{t}$ is a guesser. Recall that $L_{t}$ and $L_{t+1}$ are not missing any nodes, because $t\leq p$. Since the network is connected at round $t+1$, there is at least one node $v\in L_{t+1}$ that is guessable by $\vartheta^{\prime}$. Also, since no guess has ever been made in level $L_{t+1}$ (due to the way $t$ is updated in Line 14), it follows that $v$ is not counted and not locked. However, the algorithm cannot return $n^{\prime}=-2$ as long as there are nodes such as $v$ (Line 12). Thus, $t=p$, which means that $\vartheta^{\prime}=\vartheta^{\prime\prime}$, and hence $\vartheta^{\prime}$ has at least two children in $\mathcal{V}$, as desired.
Assume now that $p\geq s+(\ell+1)(n-1)$. If $\vartheta$ is the only node in $L_{s}$, it constitutes a counting cut. In this special case, Line 13 is never reached, $n^{\prime}=-2$ is not returned at Line 25, and $t=s$. Hence, we may assume that there are nodes in $L_{s}$ other than $\vartheta$ and Line 13 is reached (so, we have $n>1$). Recall that no nodes are missing from any level up to $L_{s+(\ell+1)(n-1)}$. In particular, no nodes are missing from the levels in the non-empty interval $\mathcal{L}$ consisting of the $(\ell+1)(n-1)$ levels from $L_{s+1}$ to $L_{s+(\ell+1)(n-1)}$. Thus, by definition of $p$, as long as no guesses are made outside of $\mathcal{L}$, 1 holds, and therefore Lemmas 5.7 and 5.8 apply. Hence, as long as no guesses are made outside of $\mathcal{L}$, there are at most $\delta(n-1)$ levels of $\mathcal{L}$ containing locked nodes (Lemma 5.7) and there are at most $n-2$ levels of $\mathcal{L}$ where all guessable nodes are already counted (Lemma 5.8). So, there are at most
@@
levels in $\mathcal{L}$ were the algorithm can make no new guesses. We conclude that $\mathcal{L}$ always contains at least one node where a guess can be made, and no guesses are ever made outside of $\mathcal{L}$ until either a counting cut is found or $n^{\prime}=-3$ is returned at Line 19. In both cases, $n^{\prime}=-2$ is not returned, and moreover $t\leq s+(\ell+1)(n-1)$, as desired.
It remains to prove that $n^{\prime}\neq-1$. Since $\vartheta$ exists by assumption, the error code $-1$ cannot be returned at Line 6 and can only be returned at Line 30. In turn, this can only occur if a counting cut has been found and $\ell^{\prime}<\ell$. However, we have already proved that no level up to $L_{t}$ is missing any nodes, which implies that the counting cut contains nodes representing all processes, and in particular $\ell^{\prime}=\delta\ell$. Since $\delta\geq 1$, we have $\ell^{\prime}\geq\ell$, and the condition at Line 30 is not satisfied. ∎
@@
Figure 8: An example of a dynamic network with $\ell=\tau=1$ and $n=6$ where the algorithm of Theorem 5.2 terminates in $33n-$ 3 rounds, almost matching the upper bound of $23n-$ 2 rounds. The white nodes in the history tree are not in the view of the leader at the last round. The numbers inside nodes represent guesses; the blue ones correspond to nodes that are marked as counted. The six highlighted nodes constitute a counting cut.

### v8

@@ 5.3 Subroutine ApproxCount
We will now define the subroutine $\texttt{ApproxCount}(\mathcal{V},s,x,\ell)$ introduced in Section 5.2 and invoked in Listing 2. We will also give a proof to the technical Lemma 5.1.
[…]
An important difficulty that is unique to the multi-leader case is that, even if the vista $\mathcal{V}$ contains a long-enough strand of leader nodes, some nodes in the strand may still be branching in the history tree, and all branches except one may be missing from $\mathcal{V}$.
@@
[…]
Conditional anonymity. The procedure ApproxCount begins by assuming that the anonymity of $\vartheta$ is $x$, and then makes deductions on other anonymities based on this assumption. Thus, for every node $v$ of $\mathcal{V}$, we distinguish between its actual anonymity $a(v)$ and its conditional anonymity, defined as $a^{\prime}(v)=\delta a(v)$. Essentially, conditional anonymities are values that ApproxCount computes under the initial assumption that $a^{\prime}(\vartheta)=x=\delta a(\vartheta)$. Clearly, these values are correct if and only if $\delta=1$.
@@
Let $u,v^{\prime}\in L_{t}$, and let $P_{1}$ and $P_{2}$ be the sets of agents represented by $u$ and $v^{\prime}$, respectively. By counting the links between $P_{1}$ and $P_{2}$ in $G_{t+1}$ in two ways, we have
@@
Our proof is by well-founded induction on $w(v)$. Assume for a contradiction that no descendants of $v$ are heavy. Observe that some descendants of $v$ have been assigned guesses, because $w(v)\geq 2$. Indeed,
$w(v)>a^{\prime}(v)=\delta a(v)\geq a(v)\geq 1.$
Thus, $v$ has $k\geq 1$ “immediate” descendants $v_{1}$, $v_{2}$, …, $v_{k}$ that have been assigned guesses. That is, for all $1\leq i\leq k$, the node $v_{i}$ has been assigned a guess, and no internal nodes of the unique black path with endpoints $v$ and $v_{i}$ have been assigned guesses.
@@
for all $1\leq i\leq k$ (recall that $a^{\prime}(v_{i})$ may not be an integer). Also,
@@
Cuts and isles. Fix a vista $\mathcal{V}$ of a history tree $\mathcal{H}$. A set of nodes $C$ in $\mathcal{V}$ is said to be a cut for a node $v\notin C$ of $\mathcal{V}$ if two conditions hold: (i) for every leaf $v^{\prime}$ of $\mathcal{V}$ that lies in the subtree hanging from $v$, the black path from $v$ to $v^{\prime}$ contains a node of $C$, and (ii) no proper subset of $C$ satisfies condition (i). A cut for the root $r$ whose nodes are all counted is said to be a counting cut.
@@
If $s$ is an isle’s root and $F$ is its set of leaves, we have $a(s)\geq\sum_{v\in F}a(v)$, which is equivalent to $a^{\prime}(s)\geq\sum_{v\in F}a^{\prime}(v)$. Note that equality does not necessarily hold, because $s$ may have some descendants in the history tree $\mathcal{H}$ that do not appear in the vista $\mathcal{V}$. If equality holds, then the isle is said to be complete. In this case, given the conditional anonymities of $s$ and of all nodes in $F$, we can easily compute the conditional anonymities of all the internal nodes by adding them up starting from the nodes in $F$ and working our way up to $s$.
Figure 7: The first levels of (a vista of) a history tree, where the colored nodes are counted. The blue nodes form a counting cut, and the orange ones define a non-trivial isle with root $s$, where the nodes with a dot are internal.
@@
1# Input: a vista $\mathcal{V}$ and three integers $s$, $x$, $\ell$
2# Output: a pair $(n^{\prime},t)$, where $n^{\prime}$ is an integer or an error message, and $t$ is an integer
@@ 32 Else return $(n^{\prime},t)$
Algorithmic details of ApproxCount. The algorithm ApproxCount uses flags to mark nodes as “guessed”, “counted”, or “locked”; initially, no node is marked. Thanks to these flags, we can check if a node $u\in\mathcal{V}$ is a guesser: let $u_{1}$, $u_{2}$, …, $u_{k}$ be the children of $u$ that are also in $\mathcal{V}$ (recall that a vista does not contain all nodes of a history tree); $u$ is a guesser if and only if it is marked as counted, all the $u_{i}$’s are marked as counted, and $a^{\prime}(u)=\sum_{i}a^{\prime}(u_{i})$ (which implies $a(u)=\sum_{i}a(u_{i})$, and thus no children of $u$ are missing from $\mathcal{V}$).
@@
The algorithm starts by assigning a conditional anonymity $a^{\prime}(\vartheta)=x$ to the first leader node $\vartheta\in L_{s}$. If no leader node exists in $L_{s}$, the error message "MissingNodes" is immediately returned (Line 6). The algorithm also finds the longest strand $P_{\vartheta}$ hanging from $\vartheta$, assigns the same conditional anonymity $x$ to all of its nodes (including the unique child of the last node of $P_{\vartheta}$) and marks them as counted (Lines 7–11). Then, as long as there are nodes that can receive a guess and no counting cut has been found, the algorithm keeps assigning guesses to nodes. A node can receive a guess if it is guessable, not counted, and not locked (Line 12).
When a guess is made on a node $v$, the node itself and all of its siblings become locked (Line 15). This is to ensure that guessed nodes will always be well spread. Moreover, as a result of $v$ becoming guessed, some nodes in the path from $v$ to its ancestor in $L_{s}$ may become heavy; if this happens, let $v^{\prime}$ be the deepest heavy node (Line 18). If $g(v^{\prime})$ is not an integer, the algorithm returns the error message "WrongGuess" (Line 19). (As we will prove later, this can only happen if $\delta\neq 1$ or some nodes in the strand $P_{\vartheta}$ have children that are not in the vista $\mathcal{V}$.) Otherwise, if $g(v^{\prime})$ is an integer, the algorithm marks $v^{\prime}$ as counted and not guessed, in accordance with Lemma 5.5. Also, since $v^{\prime}$ is no longer guessed, its siblings become unlocked and are again eligible to receive guesses (Line 20). Furthermore, if the newly counted node $v^{\prime}$ is either the root or a leaf of a complete isle $I$, then the conditional anonymities of all the internal nodes of $I$ are determined, and such nodes are marked as counted (Lines 21–24).
In the end, the algorithm performs a “reality check” and possibly returns an estimate $n^{\prime}$ of $n$ as follows. If no counting cut was found, the algorithm returns the error message "StrandTooShort" (Line 25). Otherwise, a counting cut $C$ has been found. The algorithm computes $n^{\prime}$ (respectively, $\ell^{\prime}$) as the sum of the conditional anonymities of all nodes (respectively, all leader nodes) in $C$. If $\ell^{\prime}=\ell$, then the algorithm returns $n^{\prime}$ (Line 32). Otherwise, it returns the error message "MissingNodes" if $\ell^{\prime}<\ell$ (Line 30) or the error message "WrongGuess" if $\ell^{\prime}>\ell$ (Line 31). In all cases, the algorithm also returns the maximum depth $t$ of a guessed or counted node (excluding $\vartheta$ and its descendants), or $s$ if no such node exists.
@@
Recall that a node $v$ of the history tree $\mathcal{H}$ is said to be missing from level $L_{i}$ of the vista $\mathcal{V}$ if $v$ is at the level of $\mathcal{H}$ corresponding to $L_{i}$ but does not appear in $\mathcal{V}$. Clearly, if a level of $\mathcal{V}$ has no missing nodes, all previous levels have no missing nodes, either.
@@
Observe that the subtree $\mathcal{V}^{\prime}$ of $\mathcal{V}$ formed by all levels up to $L_{t}$ satisfies all of the above conditions, as it contains $\vartheta\in L_{s}$, the root $r$, and has no counting cuts, because a counting cut for $\mathcal{V}^{\prime}$ would be a counting cut for $\mathcal{V}$, as well (recall that $\mathcal{V}$ has no counting cuts). Also, Lemma 5.6 (iii) ensures that $\mathcal{V}^{\prime}$ contains no non-trivial complete isles. However, since no nodes are missing from the levels of $\mathcal{V}^{\prime}$, all isles in $\mathcal{V}^{\prime}$ are complete, and thus must be trivial. We conclude that, if $\mathcal{V}^{\prime}$ has $f$ leaves not in the subtree hanging from $\vartheta$, it contains at most $f-1$ bad nodes. Since such leaves induce a partition of the at most $n-1$ agents not represented by $\vartheta$, we have $f\leq n-1$, implying that the number of bad nodes up to $L_{t}$ is at most $n-2$. ∎
@@
    If $n^{\prime}$ is not an error message and $t^{\prime}\geq t+n^{\prime}$ and $a(\vartheta)=a(\vartheta^{\prime})$, then $n^{\prime}=n$.
@@
[…]
@@
[…]
Let us prove statement (ii). Again, because $a(\vartheta)=a(\vartheta^{\prime})$, 1 is satisfied, and all nodes marked as counted have correct guesses. Also, $x\geq a(\vartheta)$ is equivalent to $\delta\geq 1$. By assumption, ApproxCount returns $(n^{\prime},t)$, where $n^{\prime}$ is not an error message and $t^{\prime}\geq t+n^{\prime}$. Since $n^{\prime}$ is not an error message, a counting cut $C$ was found whose nodes are within levels up to $L_{t}$, and $n^{\prime}$ is the sum of the conditional anonymities of all nodes in $C$. Let $S_{C}$ be the set of agents represented by the nodes of $C$; note that $n^{\prime}\geq|S_{C}|$, because $\delta\geq 1$. We will prove that $S_{C}$ includes all agents in the system. Assume the contrary; Lemma 3.3 implies that, since $t^{\prime}\geq t+n^{\prime}\geq t+|S_{C}|$, there is a node $z\in L_{t}$ representing some agent not in $S_{C}$. Thus, the black path from $z$ to the root $r$ does not contain any node of $C$, contradicting the fact that $C$ is a counting cut with no nodes after $L_{t}$. Therefore, $|S_{C}|=n$, i.e., the nodes in $C$ represent all agents in the system. Since ApproxCount does not return an error message, the “reality check” $\ell^{\prime}=\ell$ succeeds (Lines 30–32). However, $\ell^{\prime}$ is the sum of the conditional anonymities of all leader nodes in $C$, and hence $\ell^{\prime}=\delta\ell$, implying that $\delta=1$. Thus, $n^{\prime}=\delta n=n$, as claimed.
@@
[…]
@@
[…]
@@
Figure 8: An example of a dynamic network with $\ell=\tau=1$ and $n=6$ where the algorithm of Theorem 5.2 terminates in $33n-$ 3 rounds, almost matching the upper bound of $23n-$ 2 rounds. The white nodes in the history tree are not in the vista of the leader at the last round. The numbers inside nodes represent guesses; the blue ones correspond to nodes that are marked as counted. The six highlighted nodes constitute a counting cut.